# Improving exponential-family random graph models for bipartite networks

### v1

@@ 8 Conclusion
The existing parameters for modelling shared partners or four-cycles in ERGMs for two-mode networks frequently lead to convergence problems, especially when parameters for both modes are included in the model. A literature survey shows that, perhaps as a result, the majority of published ERGM models of two-mode networks do not include these parameters. In addition, the majority of published models do not include an assessment goodness-of-fit to four-cycles.
@@
The empirical value of the new parametrisation was demonstrated on seven empirical two-mode networks, ranging in size from hundreds of nodes to hundreds of thousands of nodes. Four of the seven networks could have models estimated with existing parameters, but in all four cases only the parameter for one of the two modes could be included. In all four cases where converged models could be estimated with the new parameters, the parameters for both modes could be included, and the fit to the four-cycle count was better than that for models using existing ERGM parameters, where these could be estimated. We failed to estimate ERGM models at all for two of the networks, and for one network a model could only be estimated using the existing parameters, and not with the new ones. The new parameters are therefore not a complete solution to the problem of estimating ERGM models of two-mode networks, but are complementary to existing parameters, and further work remains in improving the ability to fit four-cycles (and longer cycles) in bipartite networks with ERGM models.
@@
These new parameters come with both conceptual and computational costs. The conceptual cost is having to use a more general dependence class in the dependence hierarchy, without an underlying theoretical justification. The social circuit dependence assumption has theoretical justifications, as does partial inclusion ($\mathrm{PI}_{1}$) for pendant-triangle configurations [77], but we are forced to use the more general class $\mathrm{D}_{2}$ simply by the dependency induced by the new configuration, without any theoretical basis. The computational cost is incurred by having to traverse the two-neighbourhood of a dyad in computing the change statistic, and even with precomputation of two-path counts [71], this can be prohibitively slow (or require impractical amounts of memory) on dense and/or high four-cycle count or very large networks, or those with very high degree nodes, such as the Robertson pollinator network and international director interlock network.
@@
Another shortcoming is the counterintuitive and possibly confusing interpretation of the new parameters for bipartite networks (BipartiteFourCyclesNodePowerA and BipartiteFourCyclesNodePowerB; see Section 6.1). The new parameter for one-mode networks (FourCyclesNodePower) is also applicable to two-mode networks but does not treat the modes separately, and is relatively straightforward to interpret simply as a four-cycle closure parameter, so one solution is to use this, but this could omit important differences in the modes in a two mode network. Even the relatively straightforward gwdegree parameter used in statnet causes significant confusion due to the counterintuitive meaning of its sign [106, 107, 75, 108]. Having to use such relatively complex statistics, such as alternating, geometrically weighted, or the “$\alpha$-outside” or “$\alpha$-inside” [76] weightings often solves the problem of near-degeneracy, but comes at the cost of making the statistics difficult to interpret [35, p. 400]. An alternative solution to the conceptual (and possibly computational) problem is to use instead the LOLOG model [60, 59] or tapered ERGM [31, 35, 98] where a simple four-cycle parameter is unlikely to cause near-degeneracy problems, much as they enable a simple triangle parameter in one-mode networks where alternating or geometrically weighted parameters are required in standard ERGMs [59, 35, 97]. LOLOG, however, does not (currently) handle bipartite networks [59, 109], while tapered ERGM has the advantage of being implemented in the statnet framework and can use any terms in the statnet ergm package.
@@
One final potential avenue for future work is in fitting cycles of length larger than four in bipartite networks. It appears that sometimes fitting four-cycles better also fits longer cycles better, but sometimes it does not (Fig. 7). Of particular importance are six-cycles, which have been suggested as the basis (rather than four-cycles) for measuring closure in two-mode networks [10]. The dependence class required for the new statistics described in this work also admits six-cycles as configurations, however we did not attempt to fit models with six-cycles as a parameter. It seems likely that, like the simple four-cycles parameter, attempting to do so would lead to problems with near-degeneracy, necessitating the creation of another weighted configuration analogous to those defined here for four-cycles (which would then be in another, even more general, dependence class).
@@ Appendix A Models contained in literature survey
The list of 117 models in 59 publications is shown in Table LABEL:tab:literature. This table shows, for each network for which one or more models are presented in a paper, the size of both node sets in the network ($N_{A}$ is the number of nodes in node set $A$, and $N_{B}$ the number of nodes in node set $B$), the method used to estimate the model(s) in the paper, and the number of models for the network. Note that, in the case of multiple models for the same data, only a single model is counted; this is commonly the case for a model shown as developed incrementally, from a simple baseline model, and including more effects in subsequent models. In such cases, we consider only the final model. Multiple models are counted for cases such as models estimated for the same network at different time points, or for the same node set but with different edge types.
@@
Appendix B A simpler, but unsuccessful, new statistic
Wang et al. [18] define the alternating $k$-two-path statistics $\mathrm{K\text{-}C_{A}}$ and $\mathrm{K\text{-}C_{P}}$, implemented as XACA and XACB in MPNet [44, 45], and BipartiteAltKCyclesA and BipartiteAltKCyclesB in EstimNetDirected (https://github.com/stivalaa/EstimNetDirected):
$z_{\text{BipartiteAltKCyclesA}(\lambda)}=\lambda\sum_{i\in B}\sum_{\left\{l\in B\,:\,l<i\right\}}\left[1-\left(1-\frac{1}{\lambda}\right)^{L_{2}(i,l)}\right]$ (22)
where $\lambda>1$ is the decay parameter, and $L_{2}(i,l)$ is the number of two-paths connecting $i$ to $l$, as defined by equation (2) in the main text. The corresponding change statistic is [18]:
$\delta_{\text{BipartiteAltKCyclesA}(\lambda)}(i,j)$ $=\sum_{l\in B}\left[x_{il}\left(1-\frac{1}{\lambda}\right)^{L_{2}(j,l)}\right]$ (23)
$=\sum_{l\in N(i)}\left(1-\frac{1}{\lambda}\right)^{L_{2}(j,l)}.$ (24)
The statistic and change statistic for BipartiteAltKCyclesB are defined similarly.
As discussed in the main text (Section 4), one problem with this statistic, particularly if used to model bipartite closure, is that it counts two-paths (as its highest-weighted term) and not just four-cycles, and hence open paths, or cycles of length greater than four, contribute to the statistic, even if there are no four-cycles at all. A simple solution to this problem, therefore, is to remove the first term and reverse the signs, so that it no longer counts open two-paths, but the first, positive, term actually counts four-cycles. We therefore define a new statistic BipartiteAltK4CyclesA as
and its change statistic
where
is the number of two-paths connecting nodes in node set $B$ (and which therefore go through a node in node set $A$), and $\operatorname{deg}(i)$, the degree of node $i$, is the change statistic for the number of two-paths through node $i$. BipartiteAltK4CyclesB and its change statistic are defined similarly. Unlike the new FourCyclesNodePower statistics defined in the main text (Section 5), these new statistics are in the same dependence class ($\mathrm{I}_{1}$, “social circuit”) as the original $\mathrm{K\text{-}C_{A}}$ and $\mathrm{K\text{-}C_{P}}$ statistics from which they are derived.
Table B1 is a copy of Table 3 in the main text, but with the new BipartieAltK4CyclesA and BipartiteAltK4CyclesB statistics (labelled BpAK4CA and BpAK4CB respectively) included (and the $\mathrm{N_{A}}$, $\mathrm{N_{B}}$, L, BpNP4CA, and BpNP4CB columns removed to make space). This table shows the value of the new statistics on some small example networks, demonstrating, that, by design, they are zero for networks in which there are no four-cycles.
Table B1: Statistics of some example bipartite networks. C4 is the number of four-cycles,and BpAK4CA and BpAK4CB are the new statistics BipartiteAltK4CyclesA and BipartiteAltK4CyclesB, respectively. Nodes in node set A are represented as red circles, and nodes in node set B as blue squares.
Name Visualization C4 XACA XACB BpAK4CA BpAK4CB
Two-path [Uncaptioned image] 0 1 0 0 0
Four-cycle [Uncaptioned image] 1 1.5 1.5 0.5 0.5
Four-cycles-3 [Uncaptioned image] 3 4.5 1.75 1.5 1.25
Ten-cycle [Uncaptioned image] 0 5 5 0 0
Nine-star [Uncaptioned image] 0 36 0 0 0
Four-fan-3 [Uncaptioned image] 3 16.5 4.5 1.5 1.5
Unfortunately, however, simulation experiments indicate that the new BipartiteAltK4CyclesA and BipartiteAltK4CyclesB parameters are actually more problematic with respect to near-degeneracy than the original $\mathrm{K\text{-}C_{A}}$ and $\mathrm{K\text{-}C_{P}}$ parameters. Fig. B1 shows the results of simulation experiments similar to those described by Wang et al. [18, p. 19]. Bipartite networks were simulated with 30 nodes in node set $A$ and 20 nodes in node set $B$ with the Edge parameter set to $-3.0$. In three different sets of simulations, for each of the BipartiteAltKCyclesB, BipartiteAltK4CyclesB, and BipartiteFourCyclesNodePowerB parameters, the parameter in question is varied from $-1.00$ to $10.0$ in increments of $0.01$ for each of two values of $\lambda$ ($\lambda=2$ and $\lambda=5$) for BipartiteAltKCyclesB and BipartiteAltK4CyclesB, and for each of two values of $\alpha$ ($\alpha=1/2$ and $\alpha=1/5$) for BipartiteFourCyclesNodePowerB. The networks were simulated using the SimulateERGM program from the EstimNetDirected software package, using the basic ERGM sampler, with a burn-in of $10^{5}$ iterations and an interval of $10^{4}$ iterations between each of 100 samples, to ensure that samples are drawn from the equilibrium ERGM distribution, and are not too autocorrelated.
Figure B1: Effect of varying the BipartiteAltKCyclesB parameter (top), BipartiteAltK4CyclesB parameter (middle), and BipartiteFourCyclesNodePowerB parameter (bottom) on the statistic corresponding to the parameter itself (left) and the Edge statistic (right). Each graph plots the mean value of the statistic (over 100 simulations) for two different values of the relevant $\lambda$ or $\alpha$ parameter.
The top right panel in Fig. B1 shows the results of a simulation similar to that shown in Fig, 9 of Wang et al. [18]: the number of edges increases smoothly, giving good coverage of the graph space (with respect to density, at least). The top left panel shows the value of the statistic corresponding to the parameter BipartiteAltKCyclesB itself. The behaviour of this curve is starting to look as if it could be prone to near-degeneracy, with a fairly steep increase at a critical value.
The graphs in the middle row of Fig. B1 show the results for the new BipartiteAltKCyclesB parameter. In this case, both the Edge statistic (right) and BipartiteAltK4CyclesB statistic itself (left) show a phase transition, where a critical value of the parameter separates an empty graph regime from a complete graph regime. This model is therefore near-degenerate, suggesting that this new parameter may in fact be less useful than the original $\mathrm{K\text{-}C_{A}}$ and $\mathrm{K\text{-}C_{P}}$ parameters.
For completeness, the bottom row of Fig. B1 shows the results for the new BipartiteFourCyclesNodePowerB parameter defined in the main text (Section 5). When $\alpha=0.5$ an abrupt change from a near-empty to a full graph occurs, however, as shown by the result for $\alpha=0.2$, this can be removed by decreasing the value of $\alpha$.
Perhaps unsurprisingly, given the near-degeneracy demonstrated in Fig. B1, we were able to find a converged model with the new parameter for only one of the seven empirical networks described in the main text (Section 7). This is the smallest of the seven, the Inouye-Pyke pollinator web. Table B2 shows this model as Model 3 (Models 1 and 2 reproduce the models shown in Table 5 in the main text). Models with BipartiteAltK4CyclesA did not converge, and the value of $\lambda$ for BipartiteAltK4CyclesB had to be reduced from the default ($\lambda=2$) while the value of $\lambda$ for BipartiteAltStarsB had to be increased in order for this model to converge.
Table B2: ERGM parameter estimates for the Inouye-Pyke pollinator web. including Model 3 with the BipartiteAltK4Cycles parameter, with 95% confidence intervals. Estimates in bold are statistically significant at this level. “Bipartite” in effect names is abbreviated to “Bp”.
Effect Model 1 Model 2 Model 3
Edge $\mathbf{\underset{(-6.366,-5.598)}{-5.982}}$ $\mathbf{\underset{(-6.470,-5.989)}{-6.229}}$ $\mathbf{\underset{(-6.012,-5.416)}{-5.714}}$
BpAltStarsA [$\lambda=2$] $\underset{(-0.552,2.397)}{0.923}$ $\underset{(-1.674,0.878)}{-0.398}$
BpAltStarsA [$\lambda=5$] $\mathbf{\underset{(0.029,0.965)}{0.497}}$
BpAltStarsB [$\lambda=5$] $\underset{(-0.493,1.597)}{0.552}$ $\mathbf{\underset{(0.319,1.706)}{1.013}}$ $\underset{(-0.024,1.014)}{0.495}$
BpAltKCyclesB [$\lambda=2$] $\underset{(-0.322,0.332)}{0.005}$
BpFourCyclesNodePowerA [$\alpha=1/5$] $\underset{(-4.443,1.662)}{-1.391}$
BpFourCyclesNodePowerB [$\alpha=1/5$] $\mathbf{\underset{(1.670,11.401)}{6.535}}$
BpAltK4CyclesB [$\lambda=1.1$] $\mathbf{\underset{(0.004,0.102)}{0.053}}$
The interpretation of Model 3 (Table B2) is similar to that of Model 1 and Model 2 discussed in Section 7.1. The positive and significant BipartiteAltStarsA parameter indicates centralisation on pollinator nodes, that is, the presence of pollinators that pollinate many plants. The positive and significant BipartiteAltK4CyclesB parameter indicates an over-representation of pollinators (node set $A$) participating in multiple four-cycles.
Goodness-of-fit plots for Model 3 are shown in Fig B2. Compared to the goodness-of-fit for Model 2 (Fig D1), the fit to the degree distribution for mode B nodes is worse, as is the fit to the geodesic distance distribution, however the fit to the cycle length distribution is slightly better, in particular for cycles of length 12.
Figure B2: Goodness-of-fit plots for Inouye-Pyke pollinator web ERGM Model 3 (Table B2). The observed network statistics are plotted in red with the statistics of 100 simulated networks plotted as black box plots, and blue on the histograms.
In summary, we defined new BipartiteAltK4CyclesA and BipartiteAltK4CyclesB parameters as modified forms of the $\mathrm{K\text{-}C_{A}}$ and $\mathrm{K\text{-}C_{P}}$ parameters defined in Wang et al. [18], in order to count four-cycles but not open two-paths. However, near-degeneracy was exhibited in the simulation experiments (similar results occur with the larger simulated networks described in Section 6; data not shown), and a converged model using them could only be found for one of the seven networks tested. We conclude, therefore, that although these new parameters could potentially be useful in some cases, they are more prone to near-degeneracy and not as useful as either the existing $\mathrm{K\text{-}C_{A}}$ and $\mathrm{K\text{-}C_{P}}$ parameters, or the new BipartiteFourCyclesNodePowerA and BipartiteFourCyclesNodePowerB parameters described in the main text.
Appendix C Statnet implementation example
Here we demonstrate this implementation by using the the statnet ergm package [48] with the new user terms to estimate a model for the Davis “Southern Women” network [162], obtained via the latentnet R package [163, 164]. This well-known affiliation network represents the participation of 18 women (first mode) in 14 social events (second mode).
The b1np4c and b2np4c model terms take the $\alpha$ ($0<\alpha\leq 1$) value as a parameter, with a default value of $\alpha=0.5$ if omitted. For example, to estimate a model with the b2np4c (BipartiteFourCyclesNodePowerB) term with $\alpha=1/5$ (Model 4 in Table C1):
davis_model4 <- ergm(davis ~ edges + gwb1degree(1, TRUE) +
                             gwb2degree(1, TRUE) + b2np4c(1/5),
Table C1: ERGM parameter estimates for the Southern Women network, estimated with statnet. The table was generated directly from the statnet models with the texreg R package [165].
Model 1 Model 2 Model 3 Model 4
edges $-2.07^{***}$ $-0.20$ $0.47$ $-5.90^{***}$
$(0.34)$ $(0.24)$ $(0.31)$ $(0.58)$
b1star2 $0.07$
$(0.07)$
b2star2 $0.18^{***}$
$(0.04)$
gwb1deg.fixed.1 $-0.83$ $-7.04^{***}$ $10.60^{***}$
$(1.03)$ $(1.51)$ $(1.80)$
gwb2deg.fixed.1 $-2.26^{**}$ $7.76^{**}$ $-6.95^{***}$
$(0.85)$ $(2.48)$ $(1.57)$
gwb1dsp.fixed.0.5 $0.45^{***}$
$(0.12)$
gwb2dsp.fixed.0.5 $-1.33^{***}$
$(0.30)$
b2np4c.fixed.0.2 $17.30^{***}$
$(1.86)$
AIC $319.34$ $328.85$ $308.77$ $285.41$
BIC $329.93$ $339.44$ $326.41$ $299.53$
Log Likelihood $-156.67$ $-161.43$ $-149.38$ $-138.71$
${}^{***}p<0.001$; ${}^{**}p<0.01$; ${}^{*}p<0.05$
Table C1 shows four models for the Southern Women network, estimated with the stochastic approximation algorithm [24]. Model 1, with only the edges and two-star terms for each mode, is the same as [18, Model (8.3)]. Model 2, using the geometrically weighted degree terms rather than two-stars, is similar to [18, Model (8.5)], but using the statnet gwb1degree and gwb2degree terms rather than the BPNet alternating $k$-star terms $\mathrm{K\text{-}S_{P}}$ and $\mathrm{K\text{-}S_{A}}$. Note the reversal of interpretation of signs between $\mathrm{K\text{-}S_{P}}$/$\mathrm{K\text{-}S_{A}}$ and gwb1degree/gwb2degree [106, 107]. Model 3, adding the geometrically weighted dyadwise shared partner terms, is similar to [18, Model (8.6)]. Model 4 uses the new b2np4c term rather than the geometrically weighted dyadwise shared partner terms gwb1dsp and gwb2dsp (models with b1np4c did not converge).
Cycle length distribution goodness-of-fit plots for the four models are shown in Fig. C1, and statnet goodness-of-fit plots in Fig. C2. Note that all four models fit acceptably well on all the statistics included in the goodness-of-fit tests (degree distributions for each mode, dyadwise shared partners, and geodesic distance distribution), as well as the cycle length distributions. In particular, the fit to four-cycle counts is good for all models, including Model 1 and Model 2, which do not contain any terms to model four-cycles. It therefore appears that Model 1 is the most parsimonious explanation of this data, just as discussed in Wang et al. [18, pp. 22]. This model has a positive and statistically significant event two-star parameter (b2star2), indicating “greater discrepancies in the popularity of events than expected in a random network” [18, pp. 22], taking into account the density (edges) and actor two-star (b1star2) effects.
Figure C1: Cycle length distribution goodness-of-fit plots for the Southern Women network ERGM (Table C1) Model 1 (top left), Model 2 (top right), Model 3 (bottom left), and Model 4 (bottom right). Observed network statistics are plotted as red points (joined by red lines as a visual aid) with the statistics of 100 simulated networks plotted as black box plots.
Figure C2: Statnet goodness-of-fit plots for the Southern Women ERGM (Table C1) Model 1 (top left), Model 2 (top right), Model 3 (bottom left), and Model 4 (bottom right).
That ERGM models containing only terms to model density and degree distributions (and specifically not any terms to model dyadwise shared partner distributions or four-cycles) also fit the four-cycle count well indicates that the observed number of four-cycles could have occurred simply by chance [18, pp. 22]. Note that the same applies also to six-cycles. This is consistent with the results for this network described by Opsahl [10], where the observed value of the two-mode global clustering coefficient defined in that paper is not extreme in the distribution of that coefficient in random networks.
Appendix D Goodness-of-fit plots
Goodness-of-fit plots for the ERGM models for the Inouye-Pyke pollinator web (Table 5), the St Louis crime network (Table 6), the Norwegian director interlock network (Table 7), the Australian director interlock network (Table 8), the scientific collaboration network (Table 9), and the international director interlock network (Table 10) are shown in Fig. D1, Fig. D2, Fig. D3, Fig. D4, Fig. D5, and Fig. D6, respectively.
Figure D1: Goodness-of-fit plots for Inouye-Pyke pollinator web ERGM (Table 5) Model 1 (top) and Model 2 (bottom). The observed network statistics are plotted in red with the statistics of 100 simulated networks plotted as black box plots, and blue on the histograms.
Figure D2: Goodness-of-fit plots for the St Louis crime network ERGM (Table 6) Model 1 (top) and Model 2 (bottom). The observed network statistics are plotted in red with the statistics of 100 simulated networks plotted as black box plots, and blue on the histograms.
Figure D3: Goodness-of-fit plots for the Norwegian director interlock network ERGM (Table 7) Model 1 (top) and Model 2 (bottom). The observed network statistics are plotted in red with the statistics of 100 simulated networks plotted as black box plots, and blue on the histograms.
Figure D4: Goodness-of-fit plots for the Australian director interlock network ERGM (Table 8) Model 2 (top) and Model 3 (bottom). The observed network statistics are plotted in red with the statistics of 100 simulated networks plotted as black box plots, and blue on the histograms.
Figure D5: Goodness-of-fit plots for the scientific collaboration network ERGM (Table 9) Model 1 (top) and Model 2 (bottom). The observed network statistics are plotted in red with the statistics of 100 simulated networks plotted as black box plots, and blue on the histograms.
Figure D6: Goodness-of-fit plots for the international director interlock network ERGM (Table 10) Model 3. The observed network statistics are plotted in red with the statistics of 100 simulated networks plotted as black box plots, and blue on the histograms.
Appendix E Supplementary Tables
Table E1: Four-cycle descriptive statistics.
Network $\mathrm{N_{A}}$ $\mathrm{N_{B}}$ 4-cycles Unique nodes
in 4-cycles
A B
Inouye-Pyke pollinators 91 42 830 47 33
St Louis crime 870 557 144 93 121
Norwegian director interlock 1 495 367 108 89 77
Robertson pollinators 1 428 456 565 969 862 456
Australian director interlock 9 971 2 087 1 743 742 663
Scientific collaborations 16 726 22 015 70 549 7 604 14 137
International director interlock 321 869 34 769 79 885 12 567 8 136
Table E2: ERGM parameter estimates from MPNet for the Inouye-Pyke pollinator web. * indicates statistical significance at the nominal 95% level.
Effects Lambda Parameter Stderr t-ratio
XEdge 2.$0$5.5726 0.197 0.024 *
XASA 2.0000 0.7998 0.127 0.022 *
XASB 2.0000 0.7301 0.174 0.020 *
XACB 2.0000 0.0730 0.002 -0.012 *

### v3

@@ 9 Conclusion
The existing parameters for modelling shared partners or four-cycles in ERGMs for two-mode networks frequently lead to convergence problems, especially when parameters for both modes are included in the model. A literature survey shows that the majority of published ERG models of two-mode networks do not include these parameters. In addition, the majority of published models do not include an assessment goodness-of-fit to four-cycles.
@@
These new parameters come with both conceptual and computational costs. The conceptual cost is having to use a more general dependence class in the dependence hierarchy, without an underlying theoretical justification. The social circuit dependence assumption has theoretical justifications, as does partial inclusion ($\mathrm{PI}_{1}$) for pendant-triangle configurations [80], but we are forced to use the more general class $\mathrm{D}_{2}$ simply by the dependency induced by the new configuration, without any theoretical basis.
It is notable that our attempt to solve the problems with near-degeneracy observed with the existing $\mathrm{K\text{-}C_{A}}$ and $\mathrm{K\text{-}C_{A}}$ parameters with new parameters in the same dependence class ($\mathrm{I}_{1}$, “social circuit”) failed due to these parameters being even more prone to near-degeneracy (Section 5). These new parameters were motivated by the observation (Section 4) that the $\mathrm{K\text{-}C_{A}}$ and $\mathrm{K\text{-}C_{P}}$ statistics include not just four-cycles, but also simple two-paths ($k$-two-paths with $k=1$), and that this could be a cause of near-degeneracy with these parameters. However, as shown in Section 5, modifying the $\mathrm{K\text{-}C_{A}}$ and $\mathrm{K\text{-}C_{P}}$ parameters so that they only include four-cycles (and remain in the same dependency class) did not solve the problem. The problem was only solved (Section 6) by abandoning the dyadwise shared partner structure of the $\mathrm{K\text{-}C_{A}}$ and $\mathrm{K\text{-}C_{P}}$ statistics, and moving instead to a node-oriented structure, counting the number of four-cycles in which each node is involved (and using “$\alpha$-inside” weighting). This has the consequence that these new statistics are no longer in the $\mathrm{I}_{1}$ dependence class, but rather in the more general $\mathrm{D}_{2}$ class. Of course, just because we were unable to solve the problems with the $\mathrm{K\text{-}C_{A}}$ and $\mathrm{K\text{-}C_{P}}$ parameters without creating a statistic in a more general dependence class does not mean it is impossible to do so, and an interesting direction for future work could be to either try to find a statistic for four-cycles that is not prone to near-degeneracy and is in the $\mathrm{I}_{1}$ (or perhaps $\mathrm{PI}_{1}$, $\mathrm{PI}_{2}$, $\mathrm{I}_{2}$ or $\mathrm{D}_{1}$) dependence class, or to try to prove that it is not possible.
The computational cost is incurred by having to traverse the two-neighbourhood of a dyad in computing the change statistic, and even with precomputation of two-path counts [74], this can be prohibitively slow (or require impractical amounts of memory) on dense and/or high four-cycle count or very large networks, or those with very high degree nodes.
@@
Another shortcoming is the counterintuitive and possibly confusing interpretation of the new parameters for bipartite networks (BipartiteFourCyclesNodePowerA and BipartiteFourCyclesNodePowerB; see Section 6.1). The new parameter for one-mode networks (FourCyclesNodePower) is also applicable to two-mode networks but does not treat the modes separately, and is relatively straightforward to interpret simply as a four-cycle closure parameter, so one solution is to use this, but this could omit important differences in the modes in a two mode network. Even the relatively straightforward gwdegree parameter used in statnet causes significant confusion due to the counterintuitive meaning of its sign [98, 99, 78, 100]. Having to use such relatively complex statistics, such as alternating, geometrically weighted, or the “$\alpha$-outside” or “$\alpha$-inside” [79] weightings often solves the problem of near-degeneracy, but comes at the cost of making the statistics difficult to interpret [38, p. 400]. An alternative solution to the conceptual (and possibly computational) problem is to use instead the LOLOG model [63, 62] or tapered ERGM [34, 38, 101] where a simple four-cycle parameter is unlikely to cause near-degeneracy problems, much as these new models enable a simple triangle parameter in one-mode networks where alternating or geometrically weighted parameters are required in standard ERGMs [62, 38, 102]. LOLOG, however, does not (currently) handle bipartite networks [62, 103], while tapered ERGM has the advantage of being implemented in the statnet framework and can use any terms in the statnet ergm package.
@@
One final issue is fitting cycles of length larger than four in bipartite networks. Of particular importance are six-cycles, which have been suggested as the basis (rather than four-cycles) for measuring closure in two-mode networks [10, 12, 13]. The dependence class required for the new statistics described in this work also admits six-cycles as configurations, however we did not attempt to fit models with six-cycles as a parameter. It seems likely that, like the simple four-cycles parameter, attempting to do so would lead to problems with near-degeneracy, necessitating the creation of another weighted configuration analogous to those defined here for four-cycles (which would then be in another, even more general, dependence class).
@@ Appendix A Models contained in literature survey
We used Google Scholar (search date 24 November 2022) to search for papers with terms “bipartite“ and “ergm”, “bpnet” and “ergm”, “statnet” and “bipartite”, and “gwb1dsp” or “gwb2dsp” (the latter being statnet [47, 50] terms used for bipartite networks). We also included papers that we had prior knowledge of being about bipartite ERGMs, and followed backward and forward citations (using Google Scholar for the latter) to find other relevant papers.
The resulting list of 117 models in 59 publications is shown in Table LABEL:tab:literature. This table shows, for each network for which one or more models are presented in a paper, the size of both node sets in the network ($N_{A}$ is the number of nodes in node set $A$, and $N_{B}$ the number of nodes in node set $B$), the method used to estimate the model(s) in the paper, and the number of models for the network. Note that, in the case of multiple models for the same data, only a single model is counted; this is commonly the case for a model shown as developed incrementally, from a simple baseline model, and including more effects in subsequent models. In such cases, we consider only the final model. Multiple models are counted for cases such as models estimated for the same network at different time points, or for the same node set but with different edge types.
In Table 1 of the main text, the counts for the $k$-two-path parameters $\mathrm{K\text{-}C_{P}}$ and $\mathrm{K\text{-}C_{A}}$ count those estimated with $\mathrm{K\text{-}C_{P}}$ and $\mathrm{K\text{-}C_{A}}$ in BPNet, and XACA and XACB (respectively) in MPNet.
In Table 2 of the main text, the counts for gwb1dsp and gwb2dsp include those for the bipartite geometrically weighted non-edgewise shared partner terms gwb1nsp and gwb2nsp, which are equivalent to gwb1dsp and gwb2dsp, respectively, for bipartite networks.
@@
Appendix B Goodness-of-fit plots
Goodness-of-fit plots for the Southern Women network ERG models in Section 8 are shown in Fig. B1.
Figure B1: Statnet goodness-of-fit plots for the Southern Women ERGM (Table 5) Model 1 (top left), Model 2 (top right), Model 3 (bottom left), and Model 4 (bottom right).